Cane Company manufactures two products called Alpha and Beta that sell for $165 and $130, respectively. Each product uses onlRequired: 1. What is the total amount of traceable fixed manufacturing overhead for each of the two products? Alpha Beta Trac2. What is the companys total amount of common fixed expenses? Total common fixed expenses3. Assume that Cane expects to produce and sell 89,000 Alphas during the current year. One of Canes sales representatives ha5. Assume that Cane expects to produce and sell 104,000 Alphas during the current year. One of Canes sales representatives h
Answer #1

Solution 1:

Total amount of traceable fixed manufacturing overhead for the Alpha product line = 113000*$25 = $2,825,000

Total amount of traceable fixed manufacturing overhead for the Beta product line = 113000*$27 = $3,051,000

Solution 2:

Total amount of common fixed expenses = ($24 + $19) * 113000 = $4,859,000

Solution 3:

Contribution margin per unit on special order of alpha= Selling price - Variable cost per unit

= $116 - ($40 + $29 + $15 + $21) = $11 per unit

If Cane accepts the customer’s offer for 10000 alpha, then increase in operating income = 19000*$11 = $209,000

Solution 4:

Contribution margin per unit on special order of Beta= Selling price - Variable cost per unit

= $48 - ($24 + $25 + $14 + $17) = -$32

If Cane accepts the customer’s offer for 2000 Beta, then increase (decrease) in operating income = 2000*(-$32) = -$64,000

Solution 5:

Regular contribution margin per unit for Alpha = $165 - ($40 + $29 + $15 + $21) = $60 per unit

Contribution margin per unit on special order of alpha= Selling price - Variable cost per unit

= $116 - ($40 + $29 + $15 + $21) = $11 per unit

If Cane accepts the customer’s offer for 19000 alpha, additional contribution margin from special order = 19000*$11 = $209,000

If Cane accept special order then loss of contribution margin on regular order = 10000*$60 = $600,000

Incremental net operating income if the order is accepted = $209,000 - $600,000 = ($391,000)

As there is net financial disadvantage, therefore special order should not be accepted.